Find the centroid of the region bounded by y=x² and y=2


1
Expert's answer
2021-05-31T13:18:16-0400

"\text{Given} \text{ the region is bounded by }y=x^2and y=2.\newline
\text{Area},A=\int_{-\sqrt2}^{\sqrt2 }2-{x^2} dx\newline
= [2x-\frac{x^3}{3}]_{-\sqrt2}^{\sqrt2 }
\newline
=3.771\newline
\text{Let (X, Y) be the centroid.}\newline
X=\frac{\int_{-\sqrt2}^{\sqrt2 }x(2-x^2) dx}{A}\newline
=\frac{[x^2-\frac{x^4}{4}]_{-\sqrt2}^{\sqrt2 }}{3.771}\newline
=0
\newline
Y=\frac{\frac{1}{2}\int_{-\sqrt2}^{\sqrt2 }(2^2-(x^2)^2) dx}{A}\newline
=\frac{\frac{1}{2}\int_{-\sqrt2}^{\sqrt2 }(4-x^4) dx}{3.771}\newline
=\frac{[4x-\frac{x^5}{5}]_{-\sqrt2}^{\sqrt2 }}{3.771}\newline
=\frac{\frac{1}{2}×9.05}{3.771}\newline
=1.2\newline
\text{Centroid is} (0, 1.2)."